OpenStudy (warriorz13):

Omar is playing a computer game. He starts with 0 points, wins 8 points, loses 6 points 4 times, and then wins 2 points. Which integer represents Omar’s score? –30 –14 16 34

OpenStudy (warriorz13):

@noneyabusiness

OpenStudy (anonymous):

8-6=2 -6=-4 -6=-10 -6=-16 +2=-14 your answr is -14

OpenStudy (texaschic101):

0 + 8 - (6 * 4) + 2 8 - 24 + 2 10 - 24 -14 I agree with Areedy

OpenStudy (anonymous):

alright.. 0+8=8 8-6(4)=-16 -16+2=-14

OpenStudy (warriorz13):

TY

OpenStudy (texaschic101):

Great minds think alike ..lol :)

OpenStudy (anonymous):

NP. AND SO TRUE!!!